Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b.
(-3,6);8x=9y+4
The equation of the line is y=_______
(simplify the answer. By using integers or fractions for any numbers in the expression.)

Answer by mananth(12270)


-9 y = -8 x + 4 0.89
Divide by -9
y = 0.89 x + -0.44
Compare this equation with y=mx+b
slope m = 0.89

The slope of a line parallel to the above line will be the same

The slope of the required line will be 0.89

m= 0.89 ,point (-3,6)
Find b by plugging the values of m & the point in
y=mx+b
6=-2.67+b
b=8.67
m=0.89
Plug value of the slope and b
The required equation is y=0.89 x+8.67